OpenStudy (anonymous):

factor 3x2 - x - 10?

OpenStudy (saranya):

is it 3x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats what i was thinking

OpenStudy (anonymous):

x=2 or x=-5/3

OpenStudy (anonymous):

you mean (x - 2)(3x + 5)?

OpenStudy (saranya):

3x² - x - 10 = 0 =>x² - 1/6x = 10/3 + (- 1/6)² =>x² - 1/6x = 120/36 + 1/36 =>(x - 1/6)² = 121/36 =>x - 1/6 = ± 11/6 =.> x - 1/6 - 11/6, = x - 12/6, = x - 2 =>x - 1/6 + 11/6, = x + 10/6, = x + 5/3, = 3x + 5 (3x + 5)(x - 2)